# Instance Segmentation of Scene Sketches Using Natural Image Priors

### v1

@@ 3.4 Implementation Details
Figure 4: Object layers are isolated and inpainted using a pretrained SDXL model. The inpainting mask for each object is defined by intersecting overlapping masks with the object’s bounding box.
@@
Figure 5: SketchyScene dataset provides ground truth object bounding boxes and pixel-level instance segmentation masks for scene layouts.
@@
Figure 6: Samples from our synthetic dataset. We augment the SketchyScene dataset by generating vector sketches with varied drawing styles based on SketchyScene’s scene layouts. Stroke style variation is introduced by re-rendering the scenes with three different brush styles. Additionally, we create a more symbolic and challenging sketch type, resembling children’s drawings, shown on the right, while maintaining the same scene layouts.
@@
Figure 7: Illustration of our synthetic dataset. The input images are sourced from the Visual Genome dataset [13], which we filter to a subset of scenes containing 5 to 10 object instances. We generate the corresponding sketches with InstantStyle [33].
4 Scene Sketch Segmentation Benchmark
To evaluate our performance across a diverse set of sketches, we construct a synthetic annotated scene-sketch dataset. This dataset focuses on three key axes of variation, designed to extend existing datasets: (1) drawing style, (2) stroke style, and (3) object categories. We define drawing style as a spectrum ranging from symbolic, which emphasizes abstraction and simplified representation, to realistic, which prioritizes detailed and lifelike depiction. We define stroke variation as the differences in texture, width, and flow that characterize individual strokes, similar to the variety of brush types in digital drawing software. Our dataset combines two complementary pipelines to enhance diversity and object variety.
@@ 4 Scene Sketch Segmentation Benchmark
The SketchyScene dataset [47] consists of 7,265 scene layouts containing 45 object categories. These layouts are of high quality, as they were manually constructed by humans. Each data sample includes an input sketch, object class labels, bounding boxes, and a pixel-wise segmentation map (see Fig. 5). The sketches in the dataset share a consistent clipart-like style. We extend the SketchyScene dataset to include more diverse sketch styles and stroke variations. Specifically, we incorporate recent object sketching methods that introduce significantly different sketch appearances compared to SketchyScene. These include CLIPasso [29], which transforms images of individual objects into sketches with relatively high image fidelity, and SketchAgent [30], which generates symbolic sketches resembling children’s drawings, offering a more challenging out-of-distribution case. Both techniques produce vector-format sketches, which we use to assemble scene sketches while avoiding artifacts caused by transformations. Each object is placed at its ground truth location and scaled to fit its bounding box while preserving its aspect ratio. Figure 6 demonstrate sketches produced from a given SketchyScene layout. Exploring stroke variation is crucial for testing the robustness of automatic segmentation approaches, as real-world scenarios often involve highly diverse sketch styles. We augment the vector sketch using three distinct brush styles through the Adobe Illustrator Scripting API - Calligraphic Pen, Charcoal, and Brush Pen. For each brush type, we manually select the stroke width that best preserved a natural and visually appealing result.
@@
To extend the range of 45 object categories available in SketchyScene, we utilize the Visual Genome dataset [13], a large-scale dataset containing diverse and richly annotated images containing over 33,877 distinct object categories. We use InstantStyle [33], a state-of-the-art style transfer method, to generate corresponding raster sketches from the input scene images, and segment the sketch objects based on the provided image segmentation. As sketches are typically sparse, and very small objects may disappear during the translation from image to sketch, we filtered the dataset to include 1068 images containing five to ten distinct objects per scene. Our dataset expands the class categories of SketchyScene by introducing 54 additional object classes, containing in total 72 categories. A few examples of the resulting dataset are shown in Figure 7.
5 Results
Figures 1, 8, 11 and 24 present qualitative results of our method across a diverse range of sketches. These include various object categories, both abstract and detailed scenes, different styles, and sketches from our new dataset featuring stroke variations and challenging abstractions. Our method effectively handles object categories beyond those used in our fine-tuning from the SketchyScene dataset, such as toys, furniture, and food items. Our approach successfully addresses challenging scenarios, such as detailed scenes with numerous objects and occluded objects, as seen in Figure 1 and the first row of Figure 8. More results are provided in the supplementary material.
@@ 5 Results
Figure 8: Sketch segmentation results obtained by our method for a diverse set of sketch styles and levels of complexity.
Table 1: Quantitative comparisons for object detection. We report IoU, AR, and AP metrics across seven datasets, along with the mean and standard deviation for each method. IoU measures the overlap between predicted and ground-truth bounding boxes, AR evaluates the ability to detect all relevant objects, and AP combines precision and recall across varying IoU thresholds from 50% to 95%. AP@50 and AP@75 indicate Average Precision at IoU thresholds of 50% and 75%, respectively, reflecting stricter requirements for bounding box overlap. Our method demonstrates consistent improvements across nearly all datasets and metrics, significantly outperforming baselines, especially in detecting sketch objects with precision.
IoU $\uparrow$ AR $\uparrow$ AP $\uparrow$ AP@50 $\uparrow$ AP@75 $\uparrow$
@@
All $0.39$ $0.26$ $\mathbf{0.70}$ $0.29$ $0.29$ $\mathbf{0.80}$ $0.24$ $0.24$ $\mathbf{0.75}$ $0.58$ $0.34$ $\mathbf{0.88}$ $0.11$ $0.27$ $\mathbf{0.80}$
±$\pm 0.12$$\pm 0.04$$\pm 0.11$$\pm 0.08$$\pm 0.08$$\pm 0.08$$\pm 0.08$$\pm 0.07$$\pm 0.11$$\pm 0.18$$\pm 0.10$$\pm 0.08$$\pm 0.03$$\pm 0.07$$\pm 0.12$
@@
Table 2: Quantitative comparisons from image segmentation. We report Accuracy and IoU metrics across seven datasets, along with the mean and standard deviation for each method. Our method consistently outperforms baselines across all datasets.
Dataset Metric Acc $\uparrow$ IoU $\uparrow$
SketchyS G-SAM Ours SketchyS G-SAM Ours
SketchyScene $0.79$ $0.53$ $\mathbf{0.92}$ $0.72$ $0.26$ $\mathbf{0.88}$
SketchAgent $0.39$ $0.35$ $\mathbf{0.88}$ $0.38$ $0.16$ $\mathbf{0.84}$
C-Base $0.70$ $0.54$ $\mathbf{0.91}$ $0.64$ $0.32$ $\mathbf{0.88}$
C-Calligraphic $0.66$ $0.50$ $\mathbf{0.87}$ $0.63$ $0.30$ $\mathbf{0.86}$
C-Charcoal $0.59$ $0.47$ $\mathbf{0.85}$ $0.43$ $0.26$ $\mathbf{0.84}$
C-BrushPen $0.66$ $0.51$ $\mathbf{0.89}$ $0.54$ $0.29$ $\mathbf{0.85}$
InstantStyle $0.43$ $0.65$ $\mathbf{0.70}$ $0.32$ $0.44$ $\mathbf{0.78}$
All $0.60$ $0.50$ $\mathbf{0.86}$ $0.52$ $0.29$ $\mathbf{0.78}$
±$\pm 0.14$$\pm 0.09$$\pm 0.07$$\pm 0.14$$\pm 0.08$$\pm 0.03$
@@
Table 3: Quantitative comparisons from image segmentation on the filtered datasets. We report Accuracy and IoU metrics across seven datasets, along with the mean and standard deviation for each method. OpenVocab performs semantic segmentation, and requires an input text prompt. We provide ground truth class labels as input prompts to generate segmentations.
Dataset Metric Acc $\uparrow$ IoU $\uparrow$
OpenVocab Ours OpenVocab Ours
SketchyScene $0.25$ $\mathbf{0.93}$ $0.16$ $\mathbf{0.90}$
SketchAgent $0.23$ $\mathbf{0.93}$ $0.16$ $\mathbf{0.88}$
C-Base $0.29$ $\mathbf{0.94}$ $0.19$ $\mathbf{0.92}$
C-Calligraphic $0.32$ $\mathbf{0.88}$ $0.22$ $\mathbf{0.88}$
C-Charcoal $0.55$ $\mathbf{0.85}$ $0.43$ $\mathbf{0.85}$
C-BrushPen $0.26$ $\mathbf{0.91}$ $0.12$ $\mathbf{0.88}$
InstantStyle $0.26$ $\mathbf{0.79}$ $0.16$ $\mathbf{0.73}$
All $0.30$ $\mathbf{0.89}$ $0.20$ $\mathbf{0.86}$
±$\pm 0.11$$\pm 0.05$$\pm 0.10$$\pm 0.06$
5.1 Comparisons
@@
We evaluate our method alongside existing scene sketch segmentation approaches, including SketchyScene [47] and the method proposed by Bourouis et al. [2]. Additionally, we include Grounding DINO [17] as a baseline, applying it directly to sketches, using Recognize-Anything Model (RAM) [42] for automatic labeling. Our evaluation dataset consists of 7746 samples: 1,113 test samples from the SketchyScene dataset, 1,113 samples from each of the CLIPasso brush styles: Calligraphic Pen, Charcoal, and Brush Pen, 1,113 samples from the SketchAgent style, and 1068 samples from the InstantStyle sketch dataset, spanning 99 object categories. Each method was applied to these datasets following their recommended best practices. Note that the SketchyScene mask generation implementation relies on legacy dependencies that are no longer executable. Therefore, we used SAM for mask generation based on their detected bounding boxes, which yielded better performance than reported in the original paper. Figure 9 illustrates selected segmentation results for all instance segmentation methods, with additional results provided in the supplementary material. Since the method by Bourouis et al. [2] is designed for semantic segmentation rather than instance segmentation, it was evaluated separately on a filtered subset of our dataset, where each scene contained only one instance per class.
Input SketchyScene Grounded SAM Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Qualitative comparison of instance segmentation methods. Each row corresponds to a different sketch style, with the first row illustrating a sample from SketchyScene and the remaining rows showcasing samples from our dataset. Black pixels indicate regions where segmentation was not applied. Our method effectively segments sketch pixels into distinct instances without introducing artifacts, outperforming alternative approaches.
Object Detection Evaluation
For object detection, we report Intersection over Union (IoU), which measures the overlap between predicted and ground-truth bounding boxes, Average Recall (AR), which evaluates the ability to detect all relevant objects, and Average Precision (AP), which combines precision and recall across various IoU thresholds. Our goal is to assess the ability to precisely detect any object in the sketch, regardless of its class. To achieve this, we calculate the mean of these metrics across object instances rather than across classes. The results for instance segmentation methods are summarized in Tab. 1. The SketchyScene method performs well on the SketchyScene data, while its performance significantly declines across all metrics when applied to other datasets, particularly for the challenging styles of the SketchAgent samples. In contrast, our method demonstrates consistent performance across all datasets, achieving an average AR score of 0.8, as shown in the last row of the table. Notably, our method outperforms SketchyScene even on its native dataset, demonstrating the effectiveness of leveraging priors from pretrained models on natural images and adapting them for sketch segmentation. The scores obtained for our baseline method, Grounding DINO, support our claim that this model struggles to generalize to the domain of sketches without adaptation, despite its strong performance on natural images. This is evident from the large margin in scores between our method and Grounding DINO, seeing an increase of 44% in IoU, 51% increase in AR, and 51% increase in AP. Furthermore, while we fine-tuned Grounding DINO exclusively on the SketchyScene dataset, the results in the table confirm that this approach surprisingly generalizes well to very different types of sketches and object categories.
Segmentation Evaluation
We evaluate the final segmentation results using two common metrics: Pixel Accuracy (Acc), which measures the ratio of correctly labeled pixels to the total pixel count in a sketch, and Intersection over Union (IoU), which evaluates the overlap between the predicted and ground-truth segmentation masks. The results for instance segmentation methods are presented in Tab. 2, while the results for Bourouis et al. [2] are shown separately in Tab. 3 since it performs semantic segmentation and requires dataset filtering. As shown, our method outperforms alternative approaches across both metrics, with a particularly notable advantage over Grounded SAM. Additionally, our method demonstrates robustness to various styles, especially excelling on the challenging SketchAgent style compared to other methods.
6 Limitations and Future Work
While our method successfully segments scene sketches across various styles and challenging cases, it has some limitations. First, our bounding box filtering technique may still include undesired boxes, potentially introducing artifacts when combining masks into the final segmentation (Figure 10a). Second, our mask generation relies on SAM, which generally produces good masks but can occasionally introduce artifacts, particularly for objects occupying large regions in the sketch (Figure 10b). Even after applying our refinement stage, some artifacts may persist in the final segmentation. Future work could address this issue by fine-tuning SAM specifically for sketches or incorporating a learned refinement stage.
Figure 10: Examples showcasing the limitations of our approach. (a) The bounding box filtering process can still retain undesired boxes, leading to artifacts in the final segmentation. (b) SAM masks oftentimes are decent, but sometimes do contain noticeable artifacts for larger objects.
7 Conclusions
We introduced SketchSeg, a method for instance segmentation of raster scene sketches. Our approach adapts Grounding DINO, an object detection model trained on natural images, to the sketch domain through class-agnostic fine-tuning. We utilized Segment Anything (SAM) for segmentation along with a refinement stage that incorporates depth cues to resolve ambiguous pixels. Our method significantly improves upon state-of-the-art approaches in this domain, demonstrating the utility of natural image priors for sketch understanding tasks. We additionally provide a synthetic scene-level annotated sketch dataset encompassing a wide range of object categories and significant variations in drawing styles. Our experiments demonstrate that SketchSeg is robust to these variations, achieving consistent performance across diverse datasets.
Supplementary Material
[…]
Appendix A Sketch Editing Interface
Figure 13: Interactive interface for sketch editing, powered by our instance segmentation and layer completion algorithm.
Our sketch segmentation and layering technique facilitates sketch editing, allowing users to drag or manipulate segmented objects without the need to manually sketch the affected regions. We demonstrate this through an interactive sketch editing interface (Figure 13) that enables users to upload a sketch, which is then segmented and transformed into completed, ordered layers as detailed in our paper. This facilitates more efficient sketch editing by allowing artists to easily move, copy, or delete pixels associated with specific object instances, as the sketch is represented as an ordered list of layers. Please see demo video for more examples.
Appendix B Synthetic Dataset
In this section, we provide additional details on our synthetic data creation process.
B.1 Generating Vector Scene Sketches
We employ CLIPasso [29] and SketchAgent [30] to generate diverse vector sketches of single objects. For each generation method, we create 10 distinct object instances for all 45 classes in the SketchyScene dataset [47], ensuring sufficient variability in our synthetic scenes. For CLIPasso’s image-to-vector conversion, we first generate photorealistic synthetic images using SDXL [25]. The generation process uses a consistent prompt template: “A realistic image of a {class_name} with a blank background”. Figure 14 demonstrates representative pairs of synthetic input images and their corresponding generated vector sketches. For SketchAgent, we generate sketches directly from class labels as text prompts, producing 10 samples per class. Figure 15 illustrates representative examples of the generated object sketches.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Examples pairs of input synthetic image and output generated object vector sketch.
“bus” “bird” “cat” “butterfly”
Refer to caption Refer to caption Refer to caption Refer to caption
“chair” “flower” “people” “house”
Refer to caption Refer to caption Refer to caption Refer to caption
“basket” “bench” “cup” “car”
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Examples of object vector sketches generated by SketchAgent.
B.2 Generating Sketches from Natural Images
To expand beyond SketchyScene’s object categories, we employ InstantStyle [33] to transform a subset of Visual Genome [13] images into sketches, using a single CLIPasso object sketch as the style reference. Figure 16 showcases a gallery of examples from our InstantStyle-generated scene sketch dataset.
Appendix C Additional Qualitative Comparisons
We present additional qualitative comparisons between our approach and baseline methods across benchmark scene sketch datasets in Fig. 17 for SketchyScene, Fig. 18 for SketchAgent, Fig. 19 for CLIPasso, Fig. 20 for InstantStyle, to accompany our numerical evaluations included in the paper. To compare with semantic segmentation method, namely OpenVocab, by Bourouis et al. [2], we created a filtered version of all seven datasets, where each scene contains at most one instance per object class. These filtered scenes remain challenging for existing methods despite their reduced complexity. We show qualitative results in Fig.  21 for filtered SketchyScene dataset, Fig.  22 for filtered SketchAgent dataset, and Fig  23 for filtered CLIPasso dataset, to accompany our qualitative results shown in the paper.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 16: Example sketches from our InstantStyle dataset. These sketches are derived from Visual Genome  [13] containing 5 to 10 annotated objects. For visual clarity, we mask unsegmented regions in the generated sketches. This dataset contains 54 new categories beyond SketchyScene’s original 45 classes, and contain 1068 sketches in total.
Input SketchyScene Grounded SAM Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Qualitative comparison of instance segmentation methods on the SketchyScene dataset. Our method surpasses SketchyScene and Grounded SAM by delivering fine-grained, semantically consistent segmentations with precise boundaries. In the urban scene (row 1), our approach accurately segments all object instances, cleanly separating the sun, tree, building, and character from each other. For the park scenes (rows 2 and 3), it captures intricate details, such as the dog’s face and picnic items, which are either missed or over-segmented by other methods. In the residential and cottage scenes (rows 4 and 5), our method effectively delineates repetitive patterns like fences and handles dense objects like trees, preserving structural integrity where other approaches struggle. These results highlight the robustness of our method in managing complex and detailed sketches.
Input SketchyScene Grounded SAM Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: Qualitative comparison of instance segmentation methods on the SketchAgent dataset. Both SketchyScene and Grounded SAM struggle to segment these abstract sketches, which differ significantly from their training data of clipart-like and real objects. Our method successfully segments individual instances while maintaining object boundaries, even in challenging cases like the last row where objects overlap with a grid-patterned picnic blanket.
Input SketchyScene Grounded SAM Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 19: Qualitative comparison of instance segmentation methods on the CLIPasso base dataset. Our method successfully detects both large and small objects with ambiguous openings in their silhouettes, whereas baseline methods either merge multiple instances into one or fail to detect them entirely.

### v2

@@ 3.4. Implementation Details
Figure 4. Layer completion. Object layers are isolated and inpainted using a pretrained SDXL model. The inpainting mask for each object is defined by intersecting overlapping masks with the object’s bounding box.
@@
Figure 5. SketchyScene dataset provides ground truth object bounding boxes and pixel-level instance segmentation masks for scene layouts.
@@
Figure 6. Samples from our InkScenes dataset. We augment the SketchyScene dataset by generating vector sketches with varied drawing styles based on SketchyScene’s scene layouts. Stroke style variation is introduced by re-rendering the scenes with three different brush styles. Additionally, we create a more symbolic and challenging sketch type, resembling children’s drawings, shown on the right, while maintaining the same scene layouts.
@@
Figure 7. Illustration of our InkScenes dataset. The input images are sourced from the Visual Genome dataset (Krishna et al., 2017), which we filter to a subset of scenes containing 5 to 10 object instances. We generate the corresponding sketches with InstantStyle (Wang et al., 2024).
4. Scene Sketch Segmentation Benchmark
To evaluate our performance across a diverse set of sketches, we construct a synthetic annotated scene-sketch dataset: InkScenes. This dataset focuses on three key axes of variation, designed to extend existing datasets: (1) drawing style, (2) stroke style, and (3) object categories. We define drawing style as a spectrum ranging from symbolic, which emphasizes abstraction and simplified representation, to realistic, which prioritizes detailed and lifelike depiction. We define stroke variation as the differences in texture, width, and flow that characterize individual strokes, similar to the variety of brush types in digital drawing software. Our dataset combines two complementary pipelines to enhance diversity and object variety.
@@ 4. Scene Sketch Segmentation Benchmark
The SketchyScene dataset (Zou et al., 2018) consists of 7,265 scene layouts containing 45 object categories. These layouts are of high quality, as they were manually constructed by humans. Each data sample includes an input sketch, object class labels, bounding boxes, and a pixel-wise segmentation map (see Fig. 5). The sketches in the dataset share a consistent clipart-like style. We extend the SketchyScene dataset to include more diverse sketch styles and stroke variations. Specifically, we incorporate recent object sketching methods that introduce significantly different sketch appearances compared to SketchyScene. These include CLIPasso (Vinker et al., 2022), which transforms images of individual objects into sketches with relatively high image fidelity, and SketchAgent (Vinker et al., 2024), which generates symbolic sketches resembling children’s drawings, offering a more challenging out-of-distribution case. Both techniques produce vector-format sketches, which we use to assemble scene sketches while avoiding artifacts caused by transformations. Each object is placed at its ground truth location and scaled to fit its bounding box while preserving its aspect ratio. Figure 6 demonstrate sketches produced from a given SketchyScene layout. Exploring stroke variation is crucial for testing the robustness of automatic segmentation approaches, as real-world scenarios often involve highly diverse sketch styles. We augment the vector sketch using three distinct brush styles through the Adobe Illustrator Scripting API - Calligraphic Pen, Charcoal, and Brush Pen. For each brush type, we manually select the stroke width that best preserved a natural and visually appealing result.
@@
To extend the range of 45 object categories available in SketchyScene, we utilize the Visual Genome dataset (Krishna et al., 2017), a large-scale dataset containing diverse and richly annotated images containing over 33,877 distinct object categories. We use InstantStyle (Wang et al., 2024), a state-of-the-art style transfer method, to generate corresponding raster sketches from the input scene images, and segment the sketch objects based on the provided image segmentation. As sketches are typically sparse, and very small objects may disappear during the translation from image to sketch, we filtered the dataset to include 1068 images containing five to ten distinct objects per scene. Our InstantStyle subset of the InkScenes dataset includes a total of 72 categories, featuring 53 novel object classes not present in SketchyScene. A few examples of the resulting dataset are shown in Figure 7.
5. Results
Figures 1, 8, 10 and 11 present qualitative results of our method across a diverse range of sketches. These include various object categories, both abstract and detailed scenes, different styles, and sketches from our new dataset featuring stroke variations and challenging abstractions. Our method effectively handles object categories beyond those used in our fine-tuning from the SketchyScene dataset, such as toys, furniture, and food items. Our approach successfully addresses challenging scenarios, such as detailed scenes with numerous objects and occluded objects, as seen in Figure 1 and the first row of Figure 8. More results are provided in the supplementary material.
@@ 5. Results
Figure 8. InkLayer segmentation results. Our method handles sketches of diverse styles and levels of complexity. The first two rows show artist-drawn sketches spanning a range of scenarios, including wildlife scenes, crowded urban markets, and cartoon characters in various settings.
Table 1. Quantitative comparisons for object detection. We report IoU, AR, and AP metrics across eight datasets, along with the mean and standard deviation for each method. IoU measures the overlap between predicted and ground-truth bounding boxes, AR evaluates the ability to detect all relevant objects, and AP combines precision and recall across varying IoU thresholds from 50% to 95%. AP@50 and AP@75 indicate Average Precision at IoU thresholds of 50% and 75%, respectively, reflecting stricter requirements for bounding box overlap. Our method demonstrates consistent improvements across nearly all datasets and metrics, significantly outperforming baselines, especially in detecting sketch objects with precision.
Dataset Metric IoU $\uparrow$ AR $\uparrow$ AP $\uparrow$ AP@50 $\uparrow$ AP@75 $\uparrow$
@@
Zhang et al. $0.33$ $0.53$ $\mathbf{0.64}$ $0.18$ $0.48$ $\mathbf{0.69}$ $0.14$ $0.38$ $\mathbf{0.62}$ $0.31$ $0.57$ $\mathbf{0.62}$ $0.06$ $0.41$ $\mathbf{0.72}$
@@
All $0.40$ $0.32$ $\mathbf{0.70}$ $0.28$ $0.31$ $\mathbf{0.79}$ $0.23$ $0.26$ $\mathbf{0.74}$ $0.55$ $0.37$ $\mathbf{0.85}$ $0.11$ $0.29$ $\mathbf{0.79}$
±$\pm 0.10$$\pm 0.12$$\pm 0.11$$\pm 0.09$$\pm 0.10$$\pm 0.09$$\pm 0.09$$\pm 0.08$$\pm 0.12$$\pm 0.19$$\pm 0.13$$\pm 0.12$$\pm 0.04$$\pm 0.09$$\pm 0.12$
@@
Table 2. Quantitative comparisons from image segmentation. We report Accuracy and IoU metrics across eight datasets, along with the mean and standard deviation for each method. Our method consistently outperforms baselines across all datasets.
Dataset Method Acc $\uparrow$ IoU $\uparrow$
Sketchy Scene G- SAM Auto- SAM Ours Sketchy Scene G- SAM Auto- SAM Ours
SketchyScene $0.79$ $0.54$ $0.24$ $\mathbf{0.92}$ $0.72$ $0.26$ $0.17$ $\mathbf{0.88}$
Zhang et al. $0.42$ $0.72$ $0.09$ $\mathbf{0.86}$ $0.33$ $0.51$ $0.04$ $\mathbf{0.74}$
SketchAgent $0.39$ $0.35$ $0.12$ $\mathbf{0.88}$ $0.38$ $0.16$ $0.06$ $\mathbf{0.84}$
C-Base $0.70$ $0.54$ $0.08$ $\mathbf{0.91}$ $0.64$ $0.32$ $0.07$ $\mathbf{0.88}$
C-Calligraphic $0.66$ $0.50$ $0.06$ $\mathbf{0.87}$ $0.63$ $0.30$ $0.05$ $\mathbf{0.86}$
C-Charcoal $0.59$ $0.47$ $0.12$ $\mathbf{0.85}$ $0.43$ $0.26$ $0.06$ $\mathbf{0.84}$
C-BrushPen $0.66$ $0.51$ $0.05$ $\mathbf{0.89}$ $0.54$ $0.29$ $0.05$ $\mathbf{0.85}$
InstantStyle $0.43$ $0.65$ $0.25$ $\mathbf{0.70}$ $0.32$ $0.44$ $0.16$ $\mathbf{0.78}$
All $0.58$ $0.53$ $0.13$ $\mathbf{0.87}$ $0.50$ $0.32$ $0.08$ $\mathbf{0.82}$
±$\pm 0.15$$\pm 0.11$$\pm 0.08$$\pm 0.04$$\pm 0.16$$\pm 0.11$$\pm 0.05$$\pm 0.07$
@@
Table 3. Quantitative comparisons from image segmentation on the filtered datasets. We report Accuracy and IoU metrics across eight datasets, along with the mean and standard deviation for each method. Since Bourouis et al. performs semantic segmentation, and requires an input text prompt, we provide ground truth class labels as prompts to generate segmentations, and use 0.01 for confidence threshold to ensure all sketch pixels are segmented.
Dataset Method Acc $\uparrow$ IoU $\uparrow$
Bour- ouis Sketch Seger Ours Bour- ouis Sketch Seger Ours
SketchyScene $0.79$ $\mathbf{0.96}$ $0.93$ $0.58$ $\mathbf{0.90}$ $\mathbf{0.90}$
Zhang et al. $0.69$ $0.65$ $\mathbf{0.86}$ $0.50$ $0.22$ $\mathbf{0.75}$
SketchAgent $0.71$ $0.69$ $\mathbf{0.93}$ $0.52$ $0.48$ $\mathbf{0.88}$
C-Base $0.83$ $0.86$ $\mathbf{0.94}$ $0.66$ $0.73$ $\mathbf{0.92}$
C-Calligraphic $0.79$ $0.76$ $\mathbf{0.88}$ $0.61$ $0.61$ $\mathbf{0.88}$
C-Charcoal $0.80$ $0.80$ $\mathbf{0.85}$ $0.61$ $0.57$ $\mathbf{0.85}$
C-BrushPen $0.81$ $0.79$ $\mathbf{0.91}$ $0.63$ $0.59$ $\mathbf{0.88}$
InstantStyle $0.76$ $0.70$ $\mathbf{0.80}$ $0.51$ $0.45$ $\mathbf{0.75}$
All $0.77$ $0.78$ $\mathbf{0.89}$ $0.58$ $0.57$ $\mathbf{0.85}$
±$\pm 0.05$$\pm 0.10$$\pm 0.05$$\pm 0.06$$\pm 0.20$$\pm 0.07$
5.1. Comparisons
@@
We evaluate our method alongside existing scene sketch segmentation approaches, including SketchyScene (Zou et al., 2018), SketchSeger (Yang et al., 2023), and the method proposed by Bourouis et al. (2024). We include additional baselines: Grounding DINO (Liu et al., 2023b) applied directly to sketches, automatic labeling with the Recognize-Anything Model (RAM) (Zhang et al., 2023), and the Automatic SAM pipeline, which samples a 32×32 uniform grid of prompts and selects the top 50 masks to form disjoint regions. Our evaluation dataset consists of 8076 samples: $1,113$ test samples from the SketchyScene dataset, $330$ samples from the Zhang et al. dataset, $1,113$ samples from each of the CLIPasso brush styles: Base, Calligraphic Pen, Charcoal, and Brush Pen, $1,113$ samples from the SketchAgent style, and $1068$ samples from the InstantStyle sketch dataset. Each method was applied to these datasets following their recommended best practices. Note that the SketchyScene mask generation implementation relies on legacy dependencies that are no longer executable. Therefore, we used SAM for mask generation based on their detected bounding boxes, which yielded better performance than reported in the original paper. Figure 9 illustrates selected segmentation results for all instance segmentation methods, with additional results provided in the supplementary material. Since the method by Bourouis et al. (2024) and SketchSeger (2023) are designed for semantic segmentation rather than instance segmentation, it was evaluated separately on a filtered subset of our dataset, where each scene contained only one instance per class.
@@
Figure 9. Qualitative comparison of instance segmentation methods. Each row corresponds to a different sketch dataset. Black pixels indicate regions where segmentation was not applied. Our method effectively segments sketch pixels into distinct instances, outperforming alternative approaches.
Object Detection Evaluation
For object detection, we report Intersection over Union (IoU), which measures the overlap between predicted and ground-truth bounding boxes, Average Recall (AR), which evaluates the ability to detect all relevant objects, and Average Precision (AP), which combines precision and recall across various IoU thresholds. Our goal is to assess the ability to precisely detect any object in the sketch, regardless of its class. To achieve this, we calculate the mean of these metrics across object instances rather than across classes. The results for instance segmentation methods are summarized in Table 1. The SketchyScene method performs well on the SketchyScene data, while its performance significantly declines across all metrics when applied to other datasets, particularly for the challenging styles of the SketchAgent samples. In contrast, our method demonstrates consistent performance across all datasets, achieving an average AR score of 0.79, as shown in the last row of the table. Notably, our method outperforms SketchyScene even on its native dataset, demonstrating the effectiveness of leveraging priors from pretrained models on natural images and adapting them for sketch segmentation. The scores obtained for our baseline method, Grounding DINO, support our claim that this model struggles to generalize to the domain of sketches without adaptation, despite its strong performance on natural images. This is evident from the large margin in scores between our method and Grounding DINO, seeing an increase of 38% in IoU, 48% increase in AR, and 48% increase in AP. Furthermore, while we fine-tuned Grounding DINO exclusively on the SketchyScene dataset, the results in the table confirm that this approach surprisingly generalizes well to very different types of sketches and object categories.
Segmentation Evaluation
We evaluate the final segmentation results using two common metrics: Pixel Accuracy (Acc), which measures the ratio of correctly labeled pixels to the total pixel count in a sketch, and Intersection over Union (IoU), which evaluates the overlap between the predicted and ground-truth segmentation masks. The results for instance segmentation methods are presented in Table 2, while the results for Bourouis et al. (2024) and SketchSeger (2023) are shown separately in Table 3 since it performs semantic segmentation and requires dataset filtering. As shown, our method outperforms alternative approaches across both metrics, with a particularly notable advantage over Grounded SAM. While SketchSeger performs well on SketchyScene dataset, its performance degrades significantly on the symbolic style SketchAgent and struggles to generalize to novel categories included in Zhang et al. and InstantStyle styles. Meanwhile, our method demonstrates robustness to various styles, especially excelling on the challenging Zhang et al. and SketchAgent style compared to other methods.